# Wenfan Ou and Sheng Bi

### v3

@@ 2.2 Rectangularity
While the definition of RMDPs is intuitive, the tractability of MDPs cannot be straightforwardly extended to RMDPs, as it heavily relies on the rectangularity assumption. Before diving into the importance of rectangularity for RMDPs, we first deliver the definition of this strict assumption. However, various communities have studied several closely related problems at different times, and a complete consensus on a common rigorous definition has not yet been reached (Xin & Goldberg 2021). We present several popular definitions to convey the core idea of rectangularity.
@@
Lending the mathematical representations, the definition of rectangularity can be classified into three types:
@@
(1) Static form (Nilim & El Ghaoui 2005, Le Tallec 2007, Wiesemann et al. 2013). In this case, $s$-rectangular ambiguity set and $(s,a)$-rectangular ambiguity set are time-invariant. They can be represented respectively as follows:
@@
$\mathcal{P}^{\mathcal{S}}$ =$=\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\ |\ \mathbf{p}=\bigotimes_{s\in\mathcal{S}}\mathbf{p}_{s},\mathbf{p}_{s}\in\mathcal{P}_{s},\mathcal{P}_{s}\subseteq\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\right\},$
@@
$\mathcal{P}^{\mathcal{SA}}$ =$=\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\ |\ \mathbf{p}=\bigotimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbf{p}_{sa},\mathbf{p}_{sa}\in\mathcal{P}_{sa},\mathcal{P}_{sa}\subseteq\Delta^{|\mathcal{S}|}\right\}.$
@@
(2) Dynamic form (Iyengar 2005, Epstein & Schneider 2003, Iancu et al. 2015). Let $P^{+1}(\cdot|\mathcal{F}_{t})(\xi)$ be the restriction of $P(\cdot|\mathcal{F}_{t})(\xi)$ to $\mathcal{F}_{t}+1$. In this case, the ambiguity set is rectangular if it satisfies:
$\mathcal{P}_{t}(\xi)=\left\{\int P(\cdot|\mathcal{F}_{t+1})(\xi^{\prime})\mathrm{d}Q(\xi^{\prime})\ |\ P\in\mathcal{P},\forall\xi^{\prime}\in\Xi_{t+1},Q\in\mathcal{P}_{t}^{+1}(\xi)\right\},$
@@
where $\mathcal{P}_{t}^{+1}(\xi)=\left\{P^{+1}(\cdot|\mathcal{F}_{t})(\xi)\ |\ P\in\mathcal{P}\right\}$.
@@
(3) Nested form (Shapiro 2016). Let $f(\xi_{[T]})=\sum_{t=1}^{T}\gamma^{t-1}R(s_{t},a_{t},s_{t+1})+\gamma^{T}R(s_{T+1})$ for simplicity. In this case, the rectangular assumption is satisfied if the following equation holds,
$\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left[f(\xi_{[T]})\right]=\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left[\sup_{P\in\mathcal{P}}\mathbb{E}_{P|\xi_{1}}\left[\cdots\sup_{P\in\mathcal{P}}\mathbb{E}_{P|\xi_{[T-1]}}[f(\xi_{[T]})]\cdots\right]\right].$
@@
The first form is termed “static” because it lacks a subscript $t$. The static form of rectangularity aims to connect the state-wise independence with rectangularity, and a more concise representation is given by $\mathcal{P}^{\mathcal{S}}=\bigotimes_{s\in\mathcal{S}}\mathcal{P}_{s}$. It is noteworthy that the static form is often the most popular representation of rectangularity in RMDPs literature. In the risk measure community, the dynamic form of the rectangular structure is derived using one-step conditional measures, connecting rectangularity with time-(in)consistency. Recently, Shapiro (2016) proposed a nested-form condition that ensures the rectangularity assumption, directly targeting the decomposability property of the static formulation to establish the equivalence with the dynamic formulation.
While there are interesting discussions on the relationship between rectangularity and time-(in)consistency, they are beyond the scope of this review. We refer interested readers to Xin & Goldberg (2021) and the references therein for details.
@@
It is important to note that these definitions of rectangularity are not simply semantically distinct, their underlying meanings differ fundamentally as well. Though definitions differ, stage-wise independence stands as a common condition that can satisfy rectangularity as characterized in any formulation.
In the context of game theory, $(s,a)$-rectangular ambiguity sets and $s$-rectangular ambiguity sets differ in their interpretation of nature’s knowledge and commitment (Le Tallec 2007). $(s,a)$-rectangular ambiguity sets essentially assume that nature can observe the decision-maker’s actions before choosing the worst plausible realization of the transition probabilities. In other words, nature can react to the decision-maker’s actions and choose the worst-case scenario among all possible realizations of the transition probabilities that are consistent with the observed actions. This interpretation is often referred to as a reacting or adaptive nature. While $s$-rectangular ambiguity sets assume that a weaker nature must commit to a realization of the transition probabilities before observing the decision-maker’s actions. Hence, nature must choose a realization of the transition probabilities without observing the decision-maker’s actions, and this realization must be worst-case over all possible realizations that are consistent with the ambiguity sets. While different rectangular assumptions adapt to different situations, $s$-rectangular ambiguity sets derive less conservative solutions in general.
Rectangularity assumption has been investigated widely about the tractability of RMDPs, and Bagnell et al. (2001) and Wiesemann et al. (2013) prove that solving a general (non-rectangular) RMDP is NP-hard. We provide a new proof that is similar to Bagnell et al. (2001) by constructing a non-rectangular RMDP as a conjunctive normal form (CNF) formula yet more accessible. In addition, we also show how to add the rectangularity condition such that the RMDPs can be solved in polynomial time.
Proposition 2.6 (NP-hard for non-rectangularity)
@@
The CNF formula provided is used to create a non-rectangular and simplified RMDP $<\mathcal{S},\mathcal{A},\mathbf{p}>$ without discount and reward randomness. Each clause $i$ is defined as state $s_{i}$, while variable $x_{j}$ in each clause $i$ is defined as a reachable state $s_{j}$. For each $x_{j}$ in clause $i$, “True” is assigned with probability $p_{ij}$, and “False” is assigned with probability $1-p_{ij}$. Note that $p_{ij}$ is binary for a deterministic policy $\pi$. As a deterministic policy $\pi$ is selected, the worst-case scenario must be deterministic. Therefore, if “True” is assigned to variable $x_{j}$ in clause $i$, the state will move to $s_{j}$ from $s_{i}$ after taking action.
If a policy $\pi$ developed above can solve a non-rectangular robust MDP in polynomial time, it can be used to solve the CNF-SAT problem in polynomial time. However, it is well-known that solving the CNF-SAT problem is NP-hard. Therefore, we prove that solving a non-rectangular RMDP is NP-hard.
It is noteworthy that a non-rectangular RMDP may be transformed into a rectangular RMDP by introducing an additional constraint. In the general rectangular RMDPs formulation, each state is visited only once, and each visit is only to one state, making each state distinct from the others despite their possibly identical “sensation”. In the CNF setting, the variables in each clause differ from those in the other clauses, allowing for decomposition and the creation of a policy that assigns values to the variables in each clause individually. This can be solved in polynomial time, highlighting the significance of the rectangular assumption in RMDPs research.
@@
Finally, we briefly present several novel rectangular concepts to close this section. Mannor et al. (2016) propose the concept of $k$-rectangularity, which restricts the cardinality of deviations to no more than $k$ to maintain tractability. Goh et al. (2018) and Goyal & Grand-Clément (2023) decompose transition probabilities into linear combinations of $r$ factors. Essentially, they require independence of factors, instead of probabilities, and term the new assumption as $r$-rectangularity. More details about these novel rectangular concepts will be specified in Section 6. While beyond traditional rectangular assumptions, these concepts build upon the rectangularity class, which remains fundamental to RMDPs.
@@ 3 Parametric RMDPs
Parametric RMDPs are among the earliest studied RMDPs (Satia & Lave Jr 1973), which mainly assume that known parameters control transition probabilities. By predetermining parameters supports or distribution families, parametric RMDPs construct uncertainty sets in a parametric fashion. The study on parametric RMDPs has undergone significant changes, even though they fall into the same category in this paper. The development of RMDPs has seen major shifts, leading to the exploration of new approaches and techniques in parametric RMDPs.
As a pioneering work, Satia & Lave Jr (1973) propose an unprecedented formulation, known as MDPIPs, where ambiguity sets comprise individually bounded transition probabilities with given upper/lower bounds. The uncertainty sets $\mathcal{P}$ are defined by:
@@
where $p^{L}_{sas^{\prime}}$ and $p^{U}_{sas^{\prime}}$ are the predetermined lower bound and upper bound, respectively. The uncertainty set $\mathcal{P}$ in (5) builds a polyhedron. Consequently, the authors develop a policy iteration algorithm that leverages linear programming techniques in each iteration to address the challenges posed by robust formulations effectively. They prove the proposed algorithm can find an $\epsilon$-optimal policy in a finite number of iterations.
In a continuous line, Givan et al. (2000) investigate a special subclass of MDPIPs, termed as bounded-parameter MDPs (BMDPs). A BMDP $M_{\updownarrow}=(T,\mathcal{S},\mathcal{A},R_{\updownarrow},\mathbf{p}_{\updownarrow},\gamma)$ is similar with the general MDP, but $\mathbf{p}_{\updownarrow}$ and $R_{\updownarrow}$ are closed real intervals instead of real values. Compared to previous work under the general setting, the authors consider a particular MDPs family as the order-maximizing MDPs. By utilizing the special structure of this family, they show that BMDPs bound the optimal value function and directly derive the optimal solution without requiring expensive optimization techniques.
@@
More recently, Delimpaltadakis et al. (2023) also leverage the threshold structure to expand the scope of BMDPs into continuous action spaces. Distinct from simple real-valued intervals, the authors introduce two functions $f^{L},f^{U}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ to represent the lower and upper bounds on the transition probability with respect to transition $(s,a,s^{\prime})$. Due to strong duality and the ordering threshold structure, they successfully convert the original max-min formulation into a maximization problem. Furthermore, they develop an efficient value iteration algorithm based on this reformulation when $\mathcal{A}$ is a polytope
@@
Besides defining the range of transition probabilities directly, Osogami (2012) considers parametric RMDPs with $s$-rectangular uncertainty sets specified by the factor, $10<\alpha<$ 1, which determines the possibly maximal value of each transition probability based on the nominal values:
@@
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\left|0\leq p_{sas^{\prime}}\leq\frac{1}{\alpha}\hat{p}_{sas^{\prime}},\forall s^{\prime}\in\mathcal{S}\text{ and }\textstyle\sum_{s^{\prime}\in\mathcal{S}^{\prime}}p_{sas^{\prime}}=1\right\}\right..$ (6)
@@
Contrary to explicitly defining intervals for each probability, this approach just needs to determine a unique parameter $\alpha$ such that the approach can more conveniently construct ambiguity sets and achieve better generalization performance when the decision-maker has limited information about transitions. Additionally, $\alpha$ can be interpreted as the degree of robustness or risk-aware confidence level. In this paper, the author establishes the equivalence between RMDPs and risk-sensitive MDPs, where the risk confidence level is set to $1-\alpha$. This relationship provides further insights to understand and construct ambiguity sets.
Contrary to the polytopic uncertainty sets, Wiesemann et al. (2013) consider ambiguity sets $\mathcal{P}$ of the following type:
@@
$\mathcal{P}\triangleq\left.\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\right|\exists\ \theta\in\Theta\text{, such that }\mathbf{p}_{sa}:=P^{\theta}(\cdot\mid s,a),~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}\right\},$ (7)

### v4

@@ 2.2 Rectangularity
While the formulation of RMDPs is intuitive and builds on standard MDPs, some properties such as the recursive structure of MDPs do not naturally extend to RMDPs. Specifically, efficiently solving MDPs relies on the Bellman optimality equation, which decomposes the multistage problem into a recursive form. However, extending this to a robust counterpart (i.e., robust Bellman optimality equation) is not trivial, as it accounts for the global worst-case scenario, which can depend on the entire history rather than just the current state. An effective approach to ensuring the computational benefits of RMDPs is to introduce rectangular assumptions.
[…]
As rectangularity is widely applied across various fields and a complete consensus on a common rigorous definition has not yet been reached (Xin & Goldberg 2021), we present several popular definitions to convey the core idea of rectangularity.
@@
For simplicity, we term the collection of uncertainty as an ambiguity set, denoted by $\mathcal{P}$. In terms of mathematical representation, the definition of rectangularity can be categorized into three types:
@@
(1) Static form (Nilim & El Ghaoui 2005, Le Tallec 2007, Wiesemann et al. 2013). In this case, the state-wise rectangular ($s$-rectangular) ambiguity set $\mathcal{P}^{\mathcal{S}}$ and state-action-wise rectangular ($(s,a)$-rectangular) ambiguity set $\mathcal{P}^{\mathcal{SA}}$ are time-invariant. Taking the transition probability matrix, for instance, both ambiguity sets can be represented respectively as follows:
@@
$\mathcal{P}^{\mathcal{S}}$ =$=\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\,\middle|\,\mathbf{p}=\bigotimes_{s\in\mathcal{S}}\mathbf{p}_{s},\mathbf{p}_{s}\in\mathcal{P}_{s},\mathcal{P}_{s}\subseteq\Delta^{|\mathcal{S}|\times|\mathcal{A}|}\right\},$ (7)
@@
$\mathcal{P}^{\mathcal{SA}}$ =$=\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\,\middle|\,\mathbf{p}=\bigotimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbf{p}_{sa},\mathbf{p}_{sa}\in\mathcal{P}_{sa},\mathcal{P}_{sa}\subseteq\Delta^{|\mathcal{S}|}\right\}.$
where $\mathcal{P}_{s}$ and $\mathcal{P}_{sa}$ are the set of (conditional) transition probability matrices of interest.
@@
(2) Dynamic form (Iyengar 2005, Epstein & Schneider 2003, Iancu et al. 2015). This representation is closely related to the notion of time consistency. Recall $\mathbb{P}(\cdot\mid\mathcal{F}_{t})$ is the probability measure $\mathbb{P}$ conditioned on the filtration $\mathcal{F}_{t}$. Let $\mathbb{P}^{+1}(\cdot|\mathcal{F}_{t})$ be the restriction of $\mathbb{P}(\cdot\mid\mathcal{F}_{t})$ to $\mathcal{F}_{t+1}$. In this case, the ambiguity set is rectangular if it satisfies:
$\mathcal{P}_{t}=\left\{\int_{\Xi_{t+1}}\mathbb{P}(\cdot\mid\mathcal{F}_{t+1})\mathrm{d}\mathbb{Q}\,\middle|\,\mathbb{P}\in\mathcal{P},\mathbb{Q}\in\mathcal{P}_{t}^{+1}\right\},$ (8)
@@
where $\mathcal{P}_{t}^{+1}=\left\{\mathbb{P}^{+1}(\cdot\mid\mathcal{F}_{t})\mid\mathbb{P}\in\mathcal{P}\right\}$ contains the conditional one-step-ahead measures.
@@
(3) Nested form (Shapiro 2016). Let $f(\xi_{[T]})$ be the cumulative discounted reward for simplicity, where $\xi_{[T]}=(\xi_{1},\xi_{2},...,\xi_{T})$ is the history of the data up to time $T$. In the clear context, we can bridge the equivalence between $\xi$ and transition $(s,a,s^{\prime})$. In this case, the rectangular assumption is satisfied if the following equation holds,
$\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}\left[f(\xi_{[T]})\right]=\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}\left[\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}|\xi_{1}}\left[\cdots\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}|\xi_{[T-1]}}[f(\xi_{[T]})]\right]\right],$ (9)
where $\mathbb{E}_{\mathbb{P}|\xi_{[t]}}[f(\xi_{[T]})]$ is the conditional expectation of $f(\xi_{[T]})$ with respect to $\mathbb{P}\in\mathcal{P}$ given $\xi_{[t]}$.
@@
The first form is termed “static” because it lacks a subscript $t$. The static form of rectangularity aims to capture the state-wise or state-action-wise independence with rectangularity, and a more concise representation is given by $\mathcal{P}^{\mathcal{S}}=\bigotimes_{s\in\mathcal{S}}\mathcal{P}_{s}$ and $\mathcal{P}^{\mathcal{SA}}=\bigotimes_{s\in\mathcal{S},a\in\mathcal{A}}\mathcal{P}_{sa}$. It is noteworthy that the static form is often the most popular representation of rectangularity in RMDP literature. In economics, particularly risk measure community, the dynamic form of the rectangular structure is derived using one-step conditional measures, motivated by the dynamic consistency of the expected utility model. From a scenario tree perspective, this dynamic representation means that we can only consider the leaves that are from specific parent nodes. Recently, Shapiro (2016) proposed a nested-form condition that ensures the rectangularity assumption, directly targeting the decomposability property of the static formulation to establish the equivalence with the dynamic formulation. Obviously, as the supremum operator is taken over time, the right-hand side of (9) is greater than or equal to the left side. With rectangular assumption, the optimization problem of $t$ suffices to take the maximum over the corresponding set of marginal probability measures of the observations $\xi_{[t-1]}$. While there are interesting discussions on the relationship between rectangularity and time-(in)consistency, they are beyond the scope of this review. We refer interested readers to Xin & Goldberg (2021) and the references therein for details.
@@
It is necessary to recognize that these definitions of rectangularity are not simply semantically distinct but differ fundamentally in their underlying interpretations. However, no matter which of the above three definitions, with any one of the three rectangular assumptions, a recursive representation can be derived from the original multistage problem, thereby rendering the formulation tractable. Meanwhile, stage-wise independence stands as a common condition that can satisfy rectangularity as characterized in any formulation (In RMDPs, we always assume the sets of states at different stages are mutually exclusive).
Returning to the static form of rectangularity, which is most prevalent in RMDPs, we further detail the nature and distinctions between $(s,a)$-rectangular ambiguity sets and $s$-rectangular ambiguity sets. In the context of game theory, these two types of ambiguity sets differ in their assumptions regarding nature’s knowledge and commitment (Le Tallec 2007). The $(s,a)$-rectangular ambiguity sets essentially presuppose that nature can observe the decision maker’s actions before choosing the worst plausible realization of the transition probabilities. In other words, nature can react to the decision maker’s actions and choose the worst-case scenario consistent with the observed actions, often referred to as a reactive or adaptive nature. In contrast, $s$-rectangular ambiguity sets assume that a weaker nature must commit to a realization of the transition probabilities before observing the decision-maker’s actions. Here, nature must choose a realization of the transition probabilities without the benefit of observing the decision maker’s actions, ensuring that this realization is the worst case among all possible realizations consistent with the ambiguity sets. While $s$-rectangular and $(s,a)$-rectangular assumptions are suited to different contexts (e.g., whether can obtain independent transition samples for state-action pair), $s$-rectangular ambiguity sets generally (but not always) result in less conservative solutions yet with a higher computational complexity (Wiesemann et al. 2013, Ho et al. 2022). The intuition behind this lies in nature’s inability to adapt to the decision-maker’s actions under $s$-rectangular ambiguity, resulting in a less pessimistic worst-case scenario than that of $(s,a)$-rectangular sets.
The rectangular assumptions have been investigated widely about the tractability of RMDPs, and Bagnell et al. (2001) and Wiesemann et al. (2013) prove that solving a general (non-rectangular) RMDP is NP-hard. We provide a new proof that is similar to Bagnell et al. (2001) by constructing a non-rectangular RMDP as a conjunctive normal form (CNF) formula yet more straightforward. In addition, we also show how to add the rectangularity assumptions such that the RMDPs can be solved in polynomial time.
Proposition 2.6 (NP-hardness for non-rectangularity)
@@
of Proposition 2.6 We prove the NP-hardness of solving non-rectangular RMDPs by reduction from the Conjunctive Normal Form Satisfiability (CNF-SAT) problem. A Boolean formula is in CNF if it is a conjunction $\wedge$ of clauses (or a single clause), where each clause is a disjunction $\vee$ of literals (or a single literal). The CNF-SAT is the decision problem of determining if there exists an assignment of truth values (“True” or “False”) to the variables such that the given formula evaluates to “True”. For instance, the formula $(x_{1}\vee\neg x_{2})\wedge(\neg x_{1}\vee x_{3})$ is in CNF, with two clauses $C_{1}=x_{1}\vee\neg x_{2},C_{2}=\neg x_{1}\vee x_{3}$, and three literals (variables) $x_{1},x_{2},x_{3}$, where $\neg$ denotes the negation. If we assign $x_{1},x_{2},x_{3}$ all as “True”, the formula evaluates to “True”, thereby providing a solution. While the problem may seem straightforward, finding a solution is NP-hard. Next, we will demonstrate how to bridge the connection between CNF-SAT and non-rectangular RMDPs.
Let a CNF formula be $\phi=C_{1}\wedge C_{2}\wedge\cdots\ C_{m}$ over variables $x_{1},x_{2},...x_{n}$. Given the CNF formula, we construct a non-rectangular RMDP $\mathcal{M}=<\mathcal{S},\mathcal{A},\mathbf{p}>$ without discounting and reward randomness as follows:
  1. 1.
    The state space $\mathcal{S}$ consists of $n+m$ states: For each clause $C_{i}$, we create a state $s_{i}$. For each variable $x_{j}$ in clause $C_{i}$, we create state $s_{j}$ which is reachable from state $s_{i}$.
  2. 2.
    The action space $\mathcal{A}$ consists of the single action of assigning truth values to variables in clauses.
  3. 3.
    The transition probability $p_{ij}$ represents the probability of assigning “True” to variable $x_{j}$ in clause $C_{j}$. Consequently, $1-p_{ij}$ is the probability of assigning “False”. Note that the action corresponds to the assignment, thus $p_{ij}$ implicitly depends on the action, consistent with the notion of transition probability $p_{sas^{\prime}}$ in RMDPs.
Consider a deterministic policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$. Under $\pi$, each $p_{ij}$ becomes binary, either 0 or 1: if “True” is assigned to variable $x_{j}$ in clause $C_{i}$, the state transitions from $s_{i}$ to $s_{j}$ with probability 1. The constructed RMDP $\mathcal{M}$ is non-rectangular because the transition probabilities $p_{ij}$ are not independent across states. Specifically, an assignment of a truth value to $x_{j}$ in one clause (state) affects all clauses containing $x_{j}$ simultaneously. This creates a coupled uncertainty structure where the choice of action in one state constrains the possible transitions in other states.
Suppose there exists an algorithm $\mathbf{A}$ that solves $\mathcal{M}$. We can use $\mathbf{A}$ to solve the original CNF-SAT problem as follows:
  1. 1.
    Construct $\mathcal{M}$ from the given CNF formula $\phi$ as described above.
  2. 2.
    Apply algorithm $\mathbf{A}$ to find an optimal policy $\pi^{*}$ for $\mathcal{M}$.
  3. 3.
    Interpret $\pi^{*}$ as a satisfied assignment for $\phi$: if $\pi^{*}(s_{i})$ transitions to $s_{j}$, set $x_{j}$ to “True” in $C_{i}$; otherwise set in to “False”.
If algorithm $\mathbf{A}$ operates in polynomial time, this procedure would enable CNF-SAT to be solved in polynomial time. However, CNF-SAT is a known NP-complete problem. Therefore, since the reduction is polynomial-time and a solution to the RMDP provides a solution to CNF-SAT, it follows that solving non-rectangular RMDPs is NP-hard.
Now, we demonstrate how to leverage the rectangular assumption that makes constructed $\mathcal{M}$ solvable in polynomial time. In rectangular RMDPs, the assumption that each state is visited only once, and each visit is to a unique state is required. In the context of CNF-SAT reduction, if this assumption and stage-wise (clause-wise) independence are satisfied, the global consistency constraint is essentially relaxed. Formally, we remove the constraint that $x_{j}^{i}=x^{k}_{j}$ for all clauses $i$ and $k$ containing variable $j$. This relaxation makes each state (clause) independent, creating a rectangular RMDP where each state can be optimized separately. However, the solution may not be valid for the original non-rectangular RMDP or CNF-SAT problem due to potential inconsistencies. This availability of such decomposition highlights the significance of the rectangular assumption in RMDPs.
@@
Finally, we briefly present several novel rectangular concepts to close this section. Mannor et al. (2016) propose the concept of $k$-rectangularity, which restricts the cardinality of deviations to no more than $k$ to maintain tractability. Additionally, Goh et al. (2018) and Goyal & Grand-Clément (2023) represent transition probabilities as linear combinations of $r$ factors and introduce a new assumption termed $r$-rectangularity. These novel rectangular concepts render RMDPs tractable without relying on traditional rectangular assumptions (e.g., state-wise independence), at the expense of representation flexibility or independence of factors. More details about these novel rectangular concepts will be specified in Section 6. Whether using traditional rectangular assumptions or exploring novel frameworks, the principle of “no free lunch” applies. Ensuring the tractability of RMDPs inherently requires the adoption of certain (potentially strong) assumptions.
@@ 3 Parametric RMDPs
Parametric RMDPs are among the earliest studied RMDPs (Satia & Lave Jr 1973), which assume that transition probabilities are controlled by known parameters. Through predetermining supports of parameters or distribution families, parametric RMDPs transfer the uncertainty sets of transition probabilities into uncertainty sets of the parameters. Although the study of parametric RMDPs has undergone significant changes, they are grouped under the same category in this paper from a high-level perspective. These major shifts in research have led to the exploration of new approaches and techniques within the field of parametric RMDPs.
As a pioneering work, Satia & Lave Jr (1973) propose an unprecedented formulation, known as MDPs with imprecisely known parameters (MDPIPs), where uncertainty sets comprise individually bounded transition probabilities with given lower and upper bounds. The uncertainty sets $\mathcal{P}$ are defined by:
@@
where $p^{L}_{sas^{\prime}}$ and $p^{U}_{sas^{\prime}}$ are the predetermined lower and upper bounds, respectively. The uncertainty set $\mathcal{P}$ in (10) builds a polyhedron. Particularly, we argue that the authors implicitly assume the $(s,a)$-rectangular assumption. They show that there exists a pure, stationary policy that is optimal from a game theory perspective because nature always can observe the decision-maker’s decision before making its decision under $(s,a)$-rectangularity. Consequently, the authors develop a modification of the policy iteration (PI) algorithm described by Howard (1960), where the objective of the policy evaluation procedure is minimizing the expected return over any policy, instead of maximizing in the original PI algorithm. They prove the proposed algorithm can find an $\epsilon$-optimal policy in a finite number of iterations. Subsequently, White III & Eldeib (1994) study the same formulation as Satia & Lave Jr (1973), but propose a new PI algorithm that leverages the LP techniques to accelerate the convergence effectively.
Following the same line of research, Givan et al. (2000) investigate a special subclass of MDPIPs, termed as bounded-parameter MDPs (BMDPs). A BMDP, denoted as $\mathcal{M}_{\updownarrow}=<T,\mathcal{S},\mathcal{A},R_{\updownarrow},\mathbf{p}_{\updownarrow},\gamma>$, is similar with the general MDP, but $\mathbf{p}_{\updownarrow}$ and $R_{\updownarrow}$ are closed real intervals instead of fixed real values. For instance, given a transition tuple $(s,a,s^{\prime})$, the probability $p_{sas^{\prime}}$ in MDPs is a real value within range $[0,1]$. However, in BMDPs, $p_{\updownarrow,sas^{\prime}}$ is expressed as a closed interval of the form $[p_{\downarrow,sas^{\prime}},p_{\uparrow,sas^{\prime}}]$ where $0\leq p_{\downarrow,sas^{\prime}}\leq p_{\uparrow,sas^{\prime}}\leq 1$ are known parameters, similar to $p^{L}_{sas^{\prime}},p^{U}_{sas^{\prime}}$ in (10). Thus, $\mathcal{M}_{\updownarrow}$ essentially represents a set of exact MDPs, where an MDP $M\in\mathcal{M}_{\updownarrow}$ if $M=<T,\mathcal{S},\mathcal{A},R^{\prime},\mathbf{p}^{\prime},\gamma>$ and $R(s,a,s^{\prime})\in R_{\updownarrow,sas^{\prime}},p_{sas^{\prime}}\in p_{\updownarrow,sas^{\prime}}$ for all $s,s^{\prime}\in\mathcal{S},a\in\mathcal{A}$.
The construction of $\mathcal{M}_{\updownarrow}$ highlights the distinction with other parametric RMDPs, as the key elements in BMDPs are different exact MDPs, rather than uncertain transition probabilities. Given a well-defined $\mathcal{M}_{\updownarrow}$, BMDPs focus on an interval value function $V_{\updownarrow}$ which is a mapping from states to closed real intervals. For any state $s\in\mathcal{S}$, $V_{\updownarrow}(s)=[V_{\downarrow}(s),V_{\uparrow}(s)]$, where $V_{\downarrow}(s)$ and $V_{\uparrow}(s)$ are some real values. Although the element number of $\mathcal{M}_{\updownarrow}$ is generally infinite, the authors show that the attention can be restricted within a particular MDPs family as the order-maximizing MDPs. For instance, given an arbitrary state permutation $\mathcal{O}$, its transition probabilities are assigned as the following threshold structure given $s,a$:
$p_{sas_{i}}=\begin{cases}p_{\uparrow,sas_{i}},\text{ if }i\leq r-1\\ p_{\downarrow,sas_{i}},\text{ if }i\geq r\\ \end{cases}$ (11)
Where $s_{i}$ denotes the $i$-th state in the permutation $\mathcal{O}$ and $r=\arg\max_{j}\sum_{i=1}^{j-1}p_{\uparrow,sas_{i}}+\sum_{i=j}^{|\mathcal{S}|}p_{\downarrow,sas_{i}}$. In other words, an order-maximizing MDP aims to assign more transition probabilities into early states in permutation $\mathcal{O}$. Building upon this class of MDPs, the authors develop interval value iteration where $V_{\downarrow}$ is obtained in terms of a permutation with increasing value and $V_{\uparrow}$ is computed by a decreasing order. By restricting the class of MDPs to the order-maximizing ones, the computational cost is significantly reduced compared to the traditional MDPIPs with a polyhedron uncertainty set.
@@
More recently, Delimpaltadakis et al. (2023) also leverage the threshold structure like (11) to expand the scope of BMDPs into continuous action spaces. Distinct from simple real-valued intervals, the authors introduce two functions $f^{L},f^{U}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ to represent the lower and upper bounds on the transition probability with respect to transition $(s,a,s^{\prime})$. Due to strong duality and the ordering threshold structure, they successfully convert the original max-min formulation into a maximization problem. Furthermore, they develop an efficient value iteration algorithm based on this reformulation when $\mathcal{A}$ is a polytope.
@@
Besides through the lens of specific ranges of individual transition probability, Osogami (2012) considers parametric RMDPs with $s$-rectangular uncertainty sets specified by a factor, $10<\alpha<$ 1, which determines the possibly maximal value $\frac{1}{\alpha}\hat{p}_{sas^{\prime}}$ of each transition probability $p_{sas^{\prime}}$ based on the nominal values $\hat{p}_{sas^{\prime}}$:
@@
$\mathcal{P}\triangleq\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\,\left|\,0\leq p_{sas^{\prime}}\leq\frac{1}{\alpha}\hat{p}_{sas^{\prime}}\text{ and }\textstyle\sum_{s^{\prime}\in\mathcal{S}}p_{sas^{\prime}}=1\ \forall s,s^{\prime}\in\mathcal{S},a\in\mathcal{A}\right\}\right..$ (12)
@@
Instead of explicit intervals for each probability, this approach requires only a single parameter $\alpha$, which makes it more convenient to construct uncertainty sets and achieve better generalization performance when the decision-maker has highly limited information about transitions. Furthermore, one can interpret $\alpha$ as the level of robustness or risk-aware confidence level. From this perspective, the author establishes the equivalence between RMDPs and risk-sensitive MDPs, where the risk confidence level is set to be $1-\alpha$. This relationship provides further insights to understand and construct uncertainty sets.
Different from the purely polytopic uncertainty sets, Wiesemann et al. (2013) consider uncertainty sets $\mathcal{P}$ where transition probabilities follow an affine function of a vector $\theta$:
@@
$\mathcal{P}\triangleq\left.\left\{\mathbf{p}\in(\Delta^{|\mathcal{S}|})^{|\mathcal{S}|\times|\mathcal{A}|}\,\right|\,\exists\ \theta\in\Theta\text{, such that }\mathbf{p}_{sa}:=P^{\theta}(\cdot\mid s,a)\ \forall(s,a)\in\mathcal{S}\times\mathcal{A}\right\},$ (13)